The brain volumes

​(cm cubedcm3​)

of 20 brains have a mean of

1087.61087.6

cm cubedcm3

and a standard deviation of

127.3127.3

cm cubedcm3.

Use the given standard deviation and the range rule of thumb to identify the limits separating values that are significantly low or significantly high. For such​ data, would a brain volume of

1322.21322.2

cm cubedcm3

be significantly​ high?

Significantly low values are

nothing

cm cubedcm3

or lower.

​(Type an integer or a decimal. Do not​ round.)

Significantly high values are

nothing

cm cubedcm3

or higher.

​(Type an integer or a decimal. Do not​ round.)

Is

1322.21322.2

cm cubedcm3

significantly​ high?

A.

Yes comma because it is below the lower limit separating value.Yes, because it is below the lower limit separating value.

B.

​Yes, because it is between the limits separating values.

C.

Yes comma because it is above the upper limit separating value.Yes, because it is above the upper limit separating value.

D.

​No, because it is above the upper limit separating value.

E.

No comma because it is between the limits separating values.No, because it is between the limits separating values.

F.

​No, because it is below the lower limit separating value
Answer #1

If we assume 3 limit, then upper limit = = 1087.6 + (3 * 127.3) = 1087.6 + 381.9 = 1469.5 and lower limit = = 1087.6 - (3 * 127.3) = 705.7.

So, significantly low values are 705.7 or lower.

Significantly high values are 1469.5 or higher.

Since, 1322.2 is within the upper and lower limit, it is not significantly high. Hence, Option (E) is the correct choice.